# Student member, IEEE

### v1

@@ Abstract
The aerodynamic coefficients of aircrafts are significantly impacted by its geometry, especially when the angle of attack (AoA) is large. In the field of aerodynamics, traditional polynomial-based parameterization uses as few parameters as possible to describe the geometry of an airfoil. However, because the 3D geometry of a wing is more complicated than the 2D airfoil, polynomial-based parameterizations have difficulty in accurately representing the entire shape of a wing in 3D space. Existing deep learning-based methods can extract massive latent neural representations for the shape of 2D airfoils or 2D slices of wings. Recent studies highlight that directly taking geometric features as inputs to the neural networks can improve the accuracy of predicted aerodynamic coefficients. Motivated by geometry theory, we propose to incorporate Riemannian geometric features for learning Coefficient of Pressure ($C_{P}$) distributions on wing surfaces. Our method calculates geometric features (Riemannian metric, connection, and curvature) and further inputs the geometric features, coordinates and flight conditions into a deep learning model to predict the $C_{P}$ distribution. Experimental results show that our method, compared to state-of-the-art Deep Attention Network (DAN), reduces the predicted mean square error (MSE) of $C_{P}$ by an average of 8.41% for the DLR-F11 aircraft test set.
@@
Riemannian space, Geometric feature, Riemannian manifold, Deep learning.
@@ 1 INTRODUCTION
The aerodynamic coefficients of an aircraft are significantly influenced by its geometric shape, especially when the angle of attack (AoA) is large. The prevalent approach for aircraft geometry description is polynomial-based parameterization, such as Parametric Section (PARSEC) [1], Class Function/Shape Function Transformation (CST) [2], B-spline curves/surfaces [3], and Bézier curves/surfaces [4], etc. These methods typically utilize a minimal number of essential parameters to represent the geometric structure of an aircraft [5]. For example, a 4th-order CST parameterization method was used to extract 8 essential parameters as features to represent the shape of NACA0012 airfoil [6]. However, using only a small number of parameters as features makes it difficult to represent the entire geometry of wings in 3D space [7]. Instead, massive feature extraction as indicated in studies like [8, 9, 10] may provide a suitable idea for the representation of wings in 3D space.
In recent years, it has been demonstrated that deep learning, as a data-driven approach, can extract massive latent neural representations as features for geometries of objects in 3D space [11, 12, 13]. In the field of aerodynamics, there is a growing trend in utilizing deep learning-based methods to extract neural representations for aircraft shapes. Yang et al. applied a Variational Autoencoder (VAE) to optimize the shape of the GA(W)-1 airfoil according to the distribution of its Coefficient of Pressure ($C_{P}$) [14]. Du et al. combined B-spline curves with a Generative Adversarial Network (GAN) to develop BSpline-GAN, enabling the generation of a wide variety of airfoils [15]. Zahn et al. utilized a Convolutional Neural Network (CNN)-based Autoencoder to extract features from a 3D wing and predict its buffet pressure loads [16]. In the above studies, the shape of an aircraft is typically organized in the form of images or 2D slices [17]. Recent studies indicate that taking massive geometric features extracted from 2D airfoil slices as the inputs of deep learning models can improve the prediction accuracy of Coefficient of Lift ($C_{L}$), Coefficient of Drag ($C_{D}$) and $C_{P}$ [18, 19, 20], which may provide a way to predict $C_{P}$ distributions on wing surfaces in 3D space.
Motivated by both differential geometry and deep learning, we propose a learning method that incorporates Riemannian geometric features for predicting $C_{P}$ distributions on wing surfaces in 3D space  [^1] The source code of our work can be found at https://github.com/huliwei123/Incorporating-Riemannian-Geometric-Features-for-Learning-CP-Distributions-on-Airplane-Wings. This method comprises two distinct submodules: a geometric feature extraction module and a multi-feature learning module. In the geometric feature extraction module, a set of self-intersection-free Bézier surfaces is used to approximate a wing surface in 3D space to construct a piecewise smooth Riemannian manifold. Subsequently, three geometric features (Riemannian metric, connection, and curvature) at a specific point as well as its 8 neighbors are calculated from the Riemannian manifold to represent the geometry of the wing. Within the multi-feature learning module, flight conditions, coordinates, and geometric features of the wing are incorporated to predict $C_{P}$.
In summary, the contributions of our work are
  1. 1.
    We use a set of self-intersection-free Bézier surface approximating the surface of a wing in 3D space to describe its geometry;
  2. 2.
    We demonstrate that geometric features (Riemannian metric, connection and curvature) can be used to represent the geometry of wings in 3D space;
  3. 3.
    We incorporate flight conditions, coordinates and geometric features of wings in 3D space into a deep learning model to predict $C_{P}$ distributions on the wing surface, which outperforms existing methods.
The remainder of this paper is organized as follows. Section 2 provides an overview of learning methods for aerodynamic coefficient predictions based on two different aircraft shape representations. In Section 3, we introduce the construction of the piecewise smooth manifold for a wing in 3D space, and the calculation of its geometric features. Furthermore, we introduce how to incorporate geometric features, coordinates and flight conditions into a deep learning model to predict $C_{P}$ distributions on the wing surface. In Section 4, a public DLR-F11 aircraft dataset [21] is used to validate the effectiveness of our method. The conclusions of our work are presented in Section 5.
@@ 2.1 Aerodynamic coefficient predictions based on aircraft image (slice) shape representations
The above studies can predict $C_{P}$ distributions for 2D airfoils or slices. However, the aerodynamic coefficient predictions based on shape representations of aircraft manifolds, as discussed below, has been shown to further improve the prediction accuracy.
Figure 1: The structure of our proposed method. The geometric features (Riemannian metric, connection, and curvature) are calculated from an aircraft wing in 3D space by geometric feature extraction submodule. Then the flight conditions, coordinates and geometric features are taken as inputs to a multi-feature submodule to predict the $C_{P}$ distribution on wing surface.
@@ 2.2 Aerodynamic coefficient predictions based on aircraft manifold shape representations
In this method, the curved shape of an aircraft is typically mapped into a low dimensional space [32, 33], and the shape representations of aircraft are expressed by calculating geometric features based on parameters defined in the low dimensional space.
Wang et al. used cubic splines to establish a mapping between the original 2D Cartesian coordinates of UIUC airfoils and 1D curvilinear coordinates [34]. Deng et al. utilized a univalent transformation to convert the geometrical information of airfoil mesh from 2D Cartesian coordinates to curvilinear coordinates [19]. The studies mentioned above utilize curvilinear coordinates as features to represent airfoil geometry. However, fundamentally, this approach is not significantly different from using Cartesian coordinates to represent geometries, because there is no further geometric feature extractions. Xiang et al. used a set of Bézier curves to transform UIUC airfoils represented by 2D coordinates into a 1D manifold, and further calculated the Riemannian metric as a geometric feature to represent the geometry of airfoils, which results in a significant reduction in the predicted mean square error (MSE) for $C_{D}$, compared to using images (slices) as inputs [18].
@@
In summary, compared with aerodynamic coefficient predictions based on aircraft image (slice) shape representations, aerodynamic coefficient predictions based on aircraft manifold shape representations can further reduce the prediction errors of aerodynamic coefficients. However, in aerodynamic coefficient predictions based on aircraft manifold shape representations, existing studies focused on shape representations for 2D airfoils, and the challenge of this method is how to extract massive geometric features for wings in 3D space.
3 METHODOLOGY
3.1 Overview
As shown in Fig.1, our method consists of two submodules: a geometric feature extraction module and a multi-feature learning module. In the geometric feature extraction module, a set of self-intersection-free Bézier surfaces is used to approximate a wing in 3D space to form a piecewise smooth manifold, and then Riemannian metric $g_{ij}$, connection $\mit\Gamma_{ij}^{k}$ and curvature $S$ are calculated as three geometric features to represent the geometry of the wing. Because the relation between wing geometry and its $C_{P}$ distribution is highly nonlinear, and the $C_{P}$ value at a specific point $P_{ab}$ ($a$ and $b$ are the subscripts of control point matrix) is related to a neighborhood around $P_{ab}$, the geometric features of 9 points ($P_{ab}$ and its 8 neighbors) are given as inputs. In the multi-feature learning module, the flight condition $\vec{\mathbf{x}}_{1}$, the coordinate points $\vec{\mathbf{x}}_{2}$ and three geometric features ($g_{ij}$, $\mit\Gamma_{ij}^{k}$ and $S$ related to 9 points) are taken as inputs to function network_1 $\sim$ function network_5, respectively. The context network takes $[\vec{\mathbf{x}}_{1},\vec{\mathbf{x}}_{2},g_{ij},\mit\Gamma_{ij}^{k},S]$ as input. The output is $C_{P}$. The multi-feature learning structure [18, 35, 36] has been shown to be more accurate in predicting aerodynamic coefficients compared to a MLP, which explains why we use the multi-feature learning structure to predict $C_{P}$. We will further test the advantages of the multi-feature learning structure in the experiments.
@@ 3.2 The geometric feature extraction module
3.2.1 Construction of the wing manifold
In this section, we use a set of self-intersection-free [18] Bézier surfaces that approximates a wing in 3D space to form a piecewise smooth manifold (we call each part of the piecewise smooth manifold a segment).
To simplify the description, we focus on a specific segment. Consider a set of wing control points in 3D space:
@@ 3.2.1 Construction of the wing manifold
where $F(u,v)$ is a Bézier surface function. Since a Bézier surface is formed by the interweaving of two Bézier curves, the Bézier surface function $F(u,v)$ is determined by two parameters $u$ and $v$, $u,v\in[0,1]$. The variables $m$ and $n$ are the degrees of the two Bézier curves, and $B_{a,m}(u)$ and $B_{b,n}(v)$ are two coefficients of the Bézier surface.
Formula (1) can be written as
$F(u,v)=\vec{B}_{u}\cdot\left[\begin{array}[]{cccc}P_{00}&P_{01}&\cdots&P_{0n}\\ P_{10}&P_{11}&\cdots&P_{1n}\\ \vdots&\vdots&\ddots&\vdots\\ P_{m0}&P_{m1}&\cdots&P_{mn}\end{array}\right]\cdot\vec{B}_{v}$ (2)
where $\vec{B}_{u}=(B_{0,m}(u),B_{1,m}(u),\cdots,B_{m,m}(u))$ and ${\vec{B}_{v}}^{\top}=(B_{0,n}(v),B_{1,n}(v),\cdots,B_{n,n}(v))$.
@@
$\large J_{F}=d(F(u,v))=\begin{bmatrix}\frac{\partial F_{x}}{\partial u}&\frac{\partial F_{x}}{\partial v}\\[6.0pt] \frac{\partial F_{y}}{\partial u}&\frac{\partial F_{y}}{\partial u}\\[6.0pt] \frac{\partial F_{z}}{\partial u}&\frac{\partial F_{z}}{\partial u}\end{bmatrix}.$ (3)
The rank of $J_{F}$ is determined by the linearity of two vectors $(\frac{\partial F_{x}}{\partial u},\frac{\partial F_{y}}{\partial u},\frac{\partial F_{z}}{\partial u})$ and $(\frac{\partial F_{x}}{\partial v},\frac{\partial F_{y}}{\partial v},\frac{\partial F_{z}}{\partial v})$. When the two vectors are linearly independent, the rank of $J_{F}$ is 2. In this case, $F(u,v)$ is an injection [37, 38] (i.e., $F(u,v)$ is a smooth immersion [39, 38]). If the function $F(u,v)$ has no self-intersections, then $F(u,v)$ is a smooth embedding in $\mathbb{R}^{3}$ [38]. Multiple smooth embeddings are concatenated together to form a piecewise smooth manifold which is denoted as $\mathcal{M}$.
3.2.2 Calculation of geometric features
(1) Riemannian metric: A Riemannian metric $g_{ij}$ defined at an arbitrary point $p_{ab}\in\mathcal{M}$, is a smooth covariant 2-tensor field representing an inner product on the tangent space of $\mathcal{M}$ (denoted as $T_{p_{ab}}\mathcal{M}$). The function $F(u,v)$ maps a parameter combination $\left\{x_{1}=u,x_{2}=v\right\}$ in $\mathbb{R}^{2}$ to a point $P_{ab}\in\mathcal{M}$, i.e., $F(x_{1},x_{2})=P_{ab}$. Therefore, the Riemannian metric can be calculated as
@@ 3.2.2 Calculation of geometric features
where $i,j=\left\{1,2\right\}$, and $\partial_{i}=\frac{\partial}{\partial x_{i}}$ denotes the direction of the partial derivative.
(2) Riemannian connection: On a Riemannian manifold $(\mathcal{M},g)$, there exists a unique connection $\nabla$ which satisfies [40]
  1. 1.
    $\nabla$ is compatible with the metric $g$;
  2. 2.
    torsion-free.
This connection is called the Levi-Civita connection [41, 42], and we use this connection in this paper.
The connection $\nabla_{X}Y$ is a derivative of a vector field $Y$ on $(\mathcal{M},g)$ in the direction of another vector field $X$, which is also known as the covariant derivative of $Y$ along $X$ [43]. Suppose that $X=v^{i}(x)\partial_{x^{i}}$ and $Y=w^{i}(x)\partial_{x^{i}}$ are two vector fields on $(\mathcal{M},g)$, then
@@
The connection coefficient $\Gamma^{k}_{ij}$, the Christoffel symbols [44], can be written as
@@
where $g^{kl}$ denotes the inverse matrix of $g_{kl}$, and $i,j,k,l=\left\{1,2\right\}$. In Formula (6), we use the Einstein summation conventions [45]. Details of the connection formulas can be found in [40].
(3) Curvature: Curvature is a measure of whether a Riemannian manifold $(\mathcal{M},g)$ is flat or curved. Curvature is defined in the form of covariant derivative by evaluation on three vector fields $X$, $Y$, $Z$:
@@
$R(X,Y)Z$ $=R(v^{i}(x)\partial_{x^{i}},w^{j}(x)\partial_{x^{j}})u^{k}(x)\partial_{x^{k}}$ (8)
$=v^{i}(x)w^{j}(x)u^{k}(x)R(\partial_{x^{i}},\partial_{x^{j}})\partial_{x^{k}}$
$=v^{i}(x)w^{j}(x)u^{k}(x)R_{ijk}^{s}\partial_{x^{s}}.$
@@
where $i,j,k,s=\left\{1,2\right\}$. The Ricci curvature and scalar curvature [47] are used to simplify the 4-tensor $R_{ijk}^{s}$:
$R_{ij}=R_{kij}^{k}$ (10)
$S=R_{i}^{i}=g^{ij}R_{ij}$ (11)
where $R_{ij}$ denotes Ricci curvature and $S$ denotes scalar curvature. Details of curvature, Ricci curvature and scalar curvature can be found in [37].
(4) Geometric features for 9 points: We take geometric features of $P_{ab}$ and its 8 neighbors as inputs for the multi-feature learning module. Fig. 2 shows how to select the 8 neighbors for $P_{ab}$. It is worth noting that all points should be on $\mathcal{M}$, and the distance $d$ between points is a hyperparameter. If the value of $d$ is relatively small, the geometric features of neighbors and $P_{ab}$ are relatively similar.
Figure 2: The selection of neighbors.

### v2

@@ Abstract
We propose to incorporate Riemannian geometric features from the geometry of aircraft wing surfaces in the prediction of coefficient of pressure ($C_{P}$) on the aircraft wing. Contrary to existing approaches that treat the wing surface as a flat object, we represent the wing as a piecewise smooth manifold and calculate a set of Riemannian geometric features (Riemannian metric, connection, and curvature) over points of the wing. Combining these features in neighborhoods of points on the wing with coordinates and flight conditions gives inputs to a deep learning model that predicts $C_{P}$ distributions. Experimental results show that the method with incorporation of Riemannian geometric features, compared to state-of-the-art Deep Attention Network (DAN), reduces the predicted mean square error (MSE) of $C_{P}$ by an average of 15.00% for the DLR-F11 aircraft test set.
@@
Deep learning, Geometric feature, Riemannian space, Riemannian manifold, Aerodynamic performance prediction.
@@ 1 INTRODUCTION
Existing deep learning methods that treat 3D objects in the real world as flat Euclidean data utilize 2D images or slices of a 3D object as inputs to learn neural representations to reflect its geometry [1, 2]. These 3D objects, such as 3D aircraft shape [3], face structure [4] and point cloud [5], etc. have intrinsic geometric properties that can be measured within the object itself without any reference to the space it may be located in [6]. However, the discretization of 2D images and slices may lose intrinsic geometric properties of 3D objects [7, 8]. Consequently, the neural representations learned from 2D images or slices of a 3D object can not represent its intrinsic geometric integrity accurately.
Recent studies have shown that non-Euclidean space can be used to accurately represent objects with intricate geometries [9]. For examples, hyperbolic space [10] and spheres [11], two type of Riemannian manifolds, were used to represent data with tree-like geometry and cyclical geometry, respectively. In aerodynamics, there is a growing trend in treating aircraft geometries as non-Euclidean data. Deng and Wang et al. show that airfoils marked by 2D coordinate points can be represented as 1D curves (non-Euclidean manifolds) embedded in $\mathbb{R}^{2}$ [12, 13]. Further more, Xiang et al. demonstrated for the first time that the Riemannian metric, calculated from airfoil manifolds, can be used to help deep learning models improve the prediction accuracy of aerodynamic coefficients [14]. However, the above studies only focus on airfoils represented by 1D curves in $\mathbb{R}^{2}$, and the aerodynamic coefficient prediction for an arbitrary point is based on the Riemannian geometric feature calculated at the point, ignoring the geometric features calculated from neighbors around the point.
Motivated by manifold theory, we propose a Riemannian geometric features-incorporated learning (RGFiL) method  [^1] The source code of our work can be found at https://github.com/huliwei123/RGFiL. and apply it to coefficient of pressure ($C_{P}$) predictions on the wing of a 3D aircraft (in this paper, a 3D aircraft is represented as a 2D curved surface embedded in $\mathbb{R}^{3}$). RGFiL constructs a piecewise smooth manifold for a given 3D aircraft, and then a set of Riemannian geometric features (Riemannian metric, connection, and curvature) at an arbitrary point as well as its 8 neighbors are all included into consideration to represent the geometry of the aircraft. In this paper, the geometric features calculated from Riemannian manifolds are called Riemannian geometric features. At last, Riemannian geometric features, flight conditions and coordinates are input to a neural network to learn neural representations for predicting $C_{P}$ on the wing of the aircraft.
We highlight the contributions of this paper as follows: 1) This paper proposes a novel deep learning model that incorporates Riemannian geometric features for predicting $C_{P}$ distributions; 2) Experimental results show that our model has better performances in predicting $C_{P}$ distributions than existing methods without geometric features, meaning that the Riemannian geometric features are well recognized and utilized by the method; 3) The method predicts $C_{P}$ value according to the geometry of small neighborhoods on the curved surface of a 3D aircraft rather than the geometry at a single point.
The remainder of this paper is organized as follows. Section 2 provides an overview of learning methods for aerodynamic coefficient predictions based on two different aircraft geometry representations. Section 3 introduces the details of RGFiL. Section 4 shows the experimental details and results of five deep learning models on DLR-F11 aircraft dataset. The conclusions are presented in Section 5.
@@ 2.1 Aerodynamic coefficient predictions based on aircraft image (slice) shape representations
The above studies predict $C_{P}$ distributions for airfoils or 2D slices of 3D aircrafts. However, the discretization of airfoils or 2D slices may lose intrinsic geometric properties of 3D aircrafts.
@@ 2.2 Aerodynamic coefficient predictions based on aircraft manifold shape representations
In this method, the curved geometry of a 3D aircraft is typically mapped into a low dimensional space [24, 25], and the geometry representations of the aircraft are expressed by calculating Riemannian geometric features based on parameters defined in the low dimensional space.
Wang et al. used cubic splines to establish a mapping between 2D Cartesian coordinates of UIUC airfoils and 1D curvilinear coordinates [26]. Deng et al. utilized a univalent transformation to convert the geometrical information of airfoil mesh from 2D Cartesian coordinates to curvilinear coordinates [12]. Although, the curvilinear coordinates are built based on manifold theory, there is no fundamental difference between using curvilinear coordinates and using Cartesian coordinates to represent geometries, because there is no further feature extractions. Xiang et al. constructed a 1D manifold for a given 2D UIUC airfoil, and further took the calculated Riemannian metric, a Riemannian geometric feature, as input to a multi-feature learning method to predict coefficient of Drag ($C_{D}$), which results in a significant reduction compared to using images (slices) as inputs [14].
@@
In summary, compared with aerodynamic coefficient predictions based on aircraft image (slice) shape representations, aerodynamic coefficient predictions based on aircraft manifold shape representations can further reduce the prediction errors of aerodynamic coefficients. In addition, Riemannian geometric features calculated from aircraft manifolds can preserve the intrinsic geometric properties of aircraft shapes theoretically. However, existing studies focused on shape representations for 2D airfoils, and the challenge of this method is how to extract geometric features for wings in 3D space.
3 Method
Figure 1: The structure of RGFiL.
In Figure. 1, RGFiL consists of two submodules: a geometric feature extraction module and a multi-feature learning module. In the geometric feature extraction module, a set of self-intersection-free Bézier surfaces is used to approximate a 3D aircraft to form a piecewise smooth manifold, and then Riemannian metric $g_{ij}$, connection $\mit\Gamma_{ij}^{k}$ and curvature $S$ are calculated as a set of Riemannian geometric features to represent the geometry of the aircraft. Because the $C_{P}$ value at an arbitrary point is related to a neighborhood around the point, the Riemannian geometric features of the point as well as its 8 neighbors are all considered. In the multi-feature learning module, the flight condition $\vec{\mathbf{x}}_{1}$, the coordinate points $\vec{\mathbf{x}}_{2}$ and Riemannian geometric features related to 9 points are taken as inputs to predict $C_{P}$ on wing of the 3D aircraft. The multi-feature learning structure involves discovering importance of neural representations, which can be used to verify the effectiveness of Riemannian geometric features [14, 27].
@@ 3.1 The Geometric Feature Extraction Module
3.1.1 Manifold Construction
We use a set of self-intersection-free [14] Bézier surfaces that approximates a 3D aircraft to form a piecewise smooth manifold (we call each part of the piecewise smooth manifold a segment).
To simplify the description, we focus on a specific segment. Consider a set of control points:
@@ 3.1.1 Manifold Construction
where $F(u,v)$ is a Bézier surface function, $u$ and $v$ $(u,v\in[0,1])$ are two parameters of $F(u,v)$, $m$ and $n$ are the degrees of $F(u,v)$, and $B_{a,m}(u)$ and $B_{b,n}(v)$ are coefficients of $F(u,v)$.
@@
$J_{F}=d(F(u,v))=\begin{bmatrix}\frac{\partial F_{x}}{\partial u}&\frac{\partial F_{x}}{\partial v}\\[6.0pt] \frac{\partial F_{y}}{\partial u}&\frac{\partial F_{y}}{\partial u}\\[6.0pt] \frac{\partial F_{z}}{\partial u}&\frac{\partial F_{z}}{\partial u}\end{bmatrix}.$ (2)
In this case, the Rank of $J_{F}$ is 2. $F(u,v)$ is an injection [28] (i.e., $F(u,v)$ is a smooth immersion [29]). If the function $F(u,v)$ has no self-intersections, then $F(u,v)$ is a smooth embedding in $\mathbb{R}^{3}$ [28]. Multiple smooth embeddings are concatenated together to form a piecewise smooth manifold $\mathcal{M}$.
3.1.2 Calculation of Riemannian Geometric Features
1) Riemannian metric: A Riemannian metric $g_{ij}$ defined at an arbitrary point $P_{ab}\in\mathcal{M}$, is a smooth covariant 2-tensor field representing an inner product on the tangent space of $\mathcal{M}$. Let $x_{1}=u,x_{2}=v$, the Riemannian metric can be calculated as
@@ 3.1.2 Calculation of Riemannian Geometric Features
where $i,j=\left\{1,2\right\}$.
2) Connection: The connection $\nabla_{X}Y$ is a derivative of a vector field $Y$ on $(\mathcal{M},g)$ in the direction of another vector field $X$ [30]. Suppose that $X=v^{i}(x)\partial_{x^{i}}$ and $Y=w^{i}(x)\partial_{x^{i}}$, then
@@
The connection coefficient $\Gamma^{k}_{ij}$, the Christoffel symbols [31], can be written as [32]
@@
where $g^{kl}$ denotes the inverse matrix of $g_{kl}$, and $i,j,k,l=\left\{1,2\right\}$.
3) Curvature: Curvature is a measure of whether a Riemannian manifold $(\mathcal{M},g)$ is flat or curved. Curvature is defined in the form of covariant derivative by evaluation on three vector fields $X$, $Y$, $Z$:
@@
$R(X,Y)Z=v^{i}(x)w^{j}(x)u^{k}(x)R_{ijk}^{s}\partial_{x^{s}}.$ (7)
@@
where $i,j,k,s=\left\{1,2\right\}$. Usually, the Ricci curvature $R_{ij}=R_{kij}^{k}$ and scalar curvature $S=g^{ij}R_{ij}$ [34] are used to simplify $R_{ijk}^{s}$.
4) Riemannian geometric features for 9 points: RGFiL takes Riemannian geometric features of $P_{ab}$ and its 8 neighbors as inputs to the multi-feature learning module. Figure. 2 shows how to select the 8 neighbors for $P_{ab}$. All points are on $\mathcal{M}$, and the distance $d$ between points is a hyperparameter. If the value of $d$ is relatively small, the Riemannian geometric features of all points are relatively similar.
Figure 2: Neighbor selection.